# Scaling the Variational Quantum Eigensolver for Dynamic Portfolio Optimization

### v1

@@ VI Conclusion and outlooks
We find that, while all combinations of ansatz and classical optimizers successfully solve XS size (6 qubit) DPO problems, larger problems demand precise fine-tuning when configuring the VQE. We observe that in these optimization problems, the DE optimizer is highly effective for the VQE algorithm, achieving convergence in most cases. Most crucially, we find that, for these cases, the most stable and consistent VQE optimization was obtained with an ansatz tailored to both the problem and QPU properties. This highlights the need to not only consider the QPU properties, such as the connectivity and the quantum gate basis, but also consider the properties of the optimization problem while designing an ansatz.
Our work has achieved the 100-qubit utility frontier using current NISQ-era QPUs on an applied, industry-focused problem. These results set a valuable benchmark for upcoming developments on quantum computing technology and techniques. We have successfully tackled a problem with intrinsic scaling difficulties from 6 to 112 qubits. We remark that our results were obtained without any quantum error mitigation technique, which we consider essential to be addressed in future works.
@@ Acknowledgments
The authors thank Á. Díaz-Fernández, J. Luis-Hita, E. Sánchez-Martínez, and J. Fraxenet for many useful inputs and valuable discussions.
@@ Appendix A Classical optimizers
In Subsection V.1, we present the results of reproducing the implementation of the VQE in Ref. [8] to solve the DPO problem. This implementation uses a CG classical optimizer alongside a cyclic ansatz circuit [64, 8]. In particular, we use the conjugate gradient optimizer found in the SciPy minimize function [50]. All hyperparameters were kept at their default values, with the exception of the maxiter parameter, which was set to 500 iterations.
@@
The combination of the cyclic ansatz and the CG optimization method did not perform very well (see Section V.1). We attribute this outcome to two primary factors: the tendency of the CG method to converge to local minima, and the noise introduced by the cyclic ansatz architecture. The latter factor plays a particularly crucial role in DPO problems of sizes S, M, L, XL, and XXL. Both factors affect the convergence of the CG, but in different ways.
@@ Appendix E Details on classical Benchmarks
The solvers Gurobi, DOCPLEX, and GEKKO utilized in this study were based on their free versions with default configurations. For the XL and XXL experiment sizes, the execution time for DOCPLEX and Gurobi was capped at 15 minutes; for undefined time limit the models failed to converge and, after this time threshold, the optimizer does not find lower optimization cost values. It is important to note that GEKKO required an initial starting point for optimization, and the results were highly dependent on this state. In this case we found that, among all possible trivial initial states, using $x_{t,a,r}=1$ for all $t$, $a$, and $r$ give the bests results. Additionally, due to computational capacity limitations (see Appendix C), it was not feasible to obtain results for sizes greater than M using the Exhaustive Search method.
@@ Appendix F Details on Tensor Networks Calculations
To visualize the adiabatic evolution process, in Fig. 14, we show the evolution of the expected value of the Hamiltonian for the XXL size. As expected, there is a gradual decrease in the expected value, which reflects that the system is approaching the ground state of the final Hamiltonian, corresponding to the optimal solution of the given problem. The process begins with an expected value of $2.20$, which corresponds to the offset value, as it represents the state of maximum entropy, and ends at $-3.20$. According to the minimum value found by Gurobi (see 2), the final value approaches this with a relative error of approximately $0.25$. It is important to note that this is the expected value, not the minimum value of the algorithm.

### v2

@@ VI Conclusion and outlooks
We find that, while all combinations of ansatz and classical optimizers successfully solve XS size (6 qubit) DPO problems, larger problems demand precise fine-tuning when configuring the VQE. We observe that in these optimization problems, the DE optimizer is highly effective for the VQE algorithm, achieving convergence in most cases. Most crucially, we find that, for these cases, the most stable and consistent VQE optimization was obtained with an ansatz tailored to both the problem and QPU properties. This highlights the need to not only consider the QPU properties, such as the connectivity and the native quantum gates, but also consider the properties of the optimization problem while designing an ansatz.
Our work has reached the 100-qubit utility frontier using current NISQ-era QPUs on an applied, industry-focused problem. These results set a valuable benchmark for upcoming developments on quantum computing technology and techniques. We have successfully tackled a problem with intrinsic scaling difficulties from 6 to 112 qubits. We remark that our results were obtained without any quantum error mitigation technique, which we consider essential to be addressed in future works.
@@ Acknowledgments
The authors thank Á. Díaz-Fernández, J. Luis-Hita, E. Sánchez-Martínez, J. Fraxanet, and P. Rivero for many useful inputs and valuable discussions.
D. A. acknowledges the partial funding of his doctoral research at the University of Deusto, within the D4K (Deusto for Knowledge) team on applied artificial intelligence and quantum computing technologies, thanks to the support of the Basque Government.
@@ Appendix A Classical optimizers
In Subsection V.1, we present the results of reproducing the implementation of the VQE in Ref. [8] to solve the DPO problem. This implementation uses a CG classical optimizer alongside a cyclic ansatz circuit [64, 8]. In particular, we use the conjugate gradient optimizer found in the SciPy minimize function [50]. All hyperparameters were kept at their default values, with the exception of the maximum number of iterations (maxiter) parameter, which was set to 500 iterations.
@@
The combination of the cyclic ansatz and the CG optimization method is not particularly effective for our problem (see Section V.1). We attribute this outcome to two primary factors: the tendency of the CG method to converge to local minima, and the noise introduced by the cyclic ansatz architecture. The latter factor plays a particularly crucial role in DPO problems of sizes S, M, L, XL, and XXL. Both factors affect the convergence of the CG, but in different ways.
@@ Appendix E Details on classical Benchmarks
The solvers Gurobi, DOCPLEX, and GEKKO utilized in this study were based on their free versions with default configurations. For the XL and XXL experiment sizes, the execution time for DOCPLEX and Gurobi was limited to 15 minutes; for undefined time limit the models failed to converge and, after this time threshold, the optimizer does not find lower (better) optimization cost values. It is important to note that GEKKO required an initial starting point for optimization, and the results were highly dependent on this state. In this case we found that, among all possible trivial initial states, using $x_{t,a,r}=1$ for all $t$, $a$, and $r$ give the bests results.
@@ Appendix F Details on Tensor Networks Calculations
To visualize the adiabatic evolution process, in Fig. 14, we show the evolution of the expected value of the Hamiltonian for the XXL size. As expected, there is a gradual decrease in the expected value, which reflects that the system is approaching the ground state of the final Hamiltonian, corresponding to the optimal solution of the given problem. The process begins with an expected value of $2.20$, which corresponds to the offset value (as it represents the state of maximum entropy) and ends at $-3.20$. According to the minimum value found by Gurobi (see 2), the final value approaches this with a relative error of approximately $0.25$. It is important to note that this is the expected value, not the minimum value of the algorithm.